# Weak expansion properties and a large deviation principle for coarse expanding conformal systems

### v2

@@ Abstract.
Coarse expanding conformal systems were introduced by P. Haïssinsky and K. M. Pilgrim to study the essential dynamical properties of certain rational maps on the Riemann sphere in complex dynamics from the point of view of Sullivan’s dictionary. In this paper, we prove that for a metric coarse expanding conformal system $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ with repellor $X$, the map $f|_{X}\colon X\rightarrow X$ is asymptotically $h$-expansive. Moreover, we show that $f|_{X}$ is not $h$-expansive if there exists at least one branch point in the repellor. As a consequence of asymptotic $h$-expansiveness, for $f|_{X}$ and each real-valued continuous potential on $X$, there exists at least one equilibrium state. For such maps, if some additional assumptions are satisfied, we can furthermore establish a level-$2$ large deviation principle for iterated preimages, followed by an equidistribution result.
@@
The authors were partially supported by NSFC Nos. 12101017, 12090010, 12090015, BJNSF No. 1214021, and Peking University Funding No. 7101303303.
@@ 1.1. Background and motivation
In complex dynamics, the first-named author studied $h$-expansiveness and asymptotic $h$-expansiveness for expanding Thurston maps [Li15]. In this paper we consider in a more general context in complex dynamics. The dynamical systems that we study in this paper are called coarse expanding conformal (CXC) systems, introduced by P. Haïssinsky and K. M. Pilgrim [HP09]. CXC systems are a class of topological systems defined in an axiomatic way, generalizing a wide range of maps, including some non-uniformly expanding rational maps on the Riemann sphere $\widehat{\mathbb{C}}$ and expanding Thurston maps without periodic critical points studied by M. Bonk and D. Meyer [BM10, BM17] dating back to W. P. Thurston’s topological characterization of postcritially finite rational maps. Other examples of metric CXC systems (see [HP10]) include
@@
    quasiregular maps on Riemannian manifolds whose iterates are uniformly quasiregular;
@@
    smooth expanding maps on smooth compact manifolds when equipped with certain distance functions.
@@
In this paper we establish the asymptotic $h$-expansiveness of each metric CXC system $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ with repellor $X$, more precisely, of the map $f|_{X}$. The motivation to investigate asymptotic $h$-expansiveness comes from the theory of thermodynamic formalism. Investigation on concepts such as topological entropy, measure-theoretic entropy, topological pressure, and measure-theoretic pressure is a main theme in thermodynamic formalism. The maximizing measures of measure-theoretic entropy and measure-theoretic pressure are known as measures of maximal entropy and equilibrium states, respectively. The existence, uniqueness, and various other properties of equilibrium states have been studied in many contexts (see for example, [Bo75, Ru89, Pr90, KH95, MauU03, BS03, Ol03, Yu03, PU10, MayU10]).
M. Misiurewicz showed that asymptotic $h$-expansiveness of a map $f$ guarantees that the measure-theoretic entropy $\mu\mapsto h_{\mu}(f)$ is upper semi-continuous [Mi76] with respect to the $*{}^{*}$ topology. It implies that for each real-valued continuous function $\phi$, the measure-theoretic pressure $\mu\mapsto P_{\mu}(f,\phi)$ is upper semi-continuous, which guarantees the existence of equilibrium states.
@@ 1.2. Main results
We also prove that if a metric CXC system has a branch point in its repellor, then it is not $h$-expansive.
@@
Let $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ be a metric CXC system. If there exists a branch point $x_{0}\in X$, then $f|_{X}$ is not $h$-expansive.
Note that each expanding Thurston map $f\colon S^{2}\rightarrow S^{2}$ without periodic critical points is a metric CXC system (if we equip the sphere with a visual metric) with all branch points in $X$. So Theorems A and B strengthen Theorem 1.1 in [Li15] in the case where $f$ has no periodic critical points. In contrast, if an expanding Thurston map has a periodic critical point, then it is not a metric CXC system (or even a topological CXC system, see Subsection 3.2), and is known to be not asymptotically $h$-expansive (see [Li15, Theorem 1.1]).
One major difference between a Thurston map and a CXC system is that the former is defined on a topological 2-sphere, while the latter considers more general topological spaces, and that the former requires the set of forward iterated images of critical points to be finite, while the latter does not require such a strong assumption. This so-called postcritical finiteness assumption is crucial in the work of M. Bonk and D. Meyer [BM10, BM17] for the construction of certain forward invariant Jordan curves that induce Markov partitions for (sufficiently high iterations of) an expanding Thurston map (see also the work of J. W. Cannon, W. J. Floyd, and W. R. Parry [CFP07]). The proof of [Li15, Theorem 1.1] depends heavily on the combinatorial structures induced by the forward invariant Jordan curves and the corresponding Markov partitions, which are not available in the general setting of CXC systems. Our strategy in this paper builds upon the construction of bouquets (see Definition 5.6) from the more flexible good open covers.
@@
Let $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ be a metric CXC system with repellor $X$. Then for each real-valued continuous function $\phi\colon X\rightarrow\mathbb{R}$, there exists at least one equilibrium state for the map $f|_{X}$ and potential $\phi$. In particular, there exists at least one measure of maximal entropy for $f|_{X}$.
Applying some additional assumptions, we furthermore establish a level-$2$ large deviation principle (see Subsection 7.1 for the definition) for metric CXC systems.
@@
satisfies a level-$2$ large deviation principle with rate function $I^{\phi}$, and it converges in the $*{}^{*}$ topology to the Dirac measure supported on the unique equilibrium state for $f$ and potential $\phi$. Furthermore, for each convex open subset $\mathfrak{G}$ of $\mathcal{P}(X)$ containing some invariant measure, we have
@@
Under proper assumptions, there exists a unique equilibrium state for the map $f|_{X}$ and each Hölder continuous potential. It follows from the level-$2$ large deviation principle that iterated preimages are equidistributed with respect to the equilibrium state.
Theorem E (Equidistribution).
  1. (1)
    $\mathfrak{X}_{0}$ and $\mathfrak{X}_{1}$ are strongly path-connected.
  2. (2)
    The branch set $B_{f}$ is finite.
Let $\phi$ be a real-valued Hölder continuous function on $X$, and $\mu_{\phi}$ be the unique equilibrium state.
Let $\{x_{n}\}_{n\in\mathbb{N}}$ be a sequence of points in $X$. Consider the following sequence of Borel probability measures on $X$:
where $S_{n}$ is defined as in Theorem D. Then $\nu_{n}$ converges to $\mu_{\phi}$ in the $*{}^{*}$ topology as $n$ tends to infinity.
Our results in this paper partially extend results in [HP09, BM10, BM17, Li15].
1.3. Structure and strategies
We will now give a brief description of the structure of this paper and our strategies.
@@ 1.3. Structure and strategies
In Section 5, we investigate the weak expansion properties of metric CXC systems.
We first introduce some basic concepts in Subsection 5.1. We review the notion of topological conditional entropy $h(g|\eta)$ of a continuous map $g\colon Z\rightarrow Z$ given an open cover $\eta$ of $Z$, and the notion of topological tail entropy $h^{*}(g)$. We then define $h$-expansiveness and asymptotic $h$-expansiveness using these notions.
In Subsection 5.2, we prove several lemmas that will be used in the proof of the asymptotic $h$-expansiveness. In particular, Lemma 5.9 gives a cover $E_{m}(U_{0}^{m},\cdots,U_{n-1}^{m};U_{m}^{n})$ of each set of the form $\bigcap_{i=0}^{n}f^{-i}{(U_{i}^{m}\cap X)}$ by level-$(m+n)$ good open sets, where $m\in\mathbb{N}_{0}$, $n\in\mathbb{N}$, and each $U_{i}^{m}\in\mathcal{U}_{m}$ is a level-$m$ good open set. In Lemma 5.10, we cover level-$m$ good open sets with level-$(m+k)$ good open sets, where $m\in\mathbb{N}_{0}$ and $k\in\mathbb{N}$.
Subsections 5.3 consists of the proofs of Theorems A and B, respectively. To prove Theorem A, for each set in the form $\bigcap_{i=0}^{n}f^{-i}(U_{i}^{m}\cap X)$, we give an upper bound for $\operatorname{card}(E_{m}(U_{0}^{m},\cdots,U_{n-1}^{m};U_{m}^{n}))$ that depends only on $m,\,n$. Combining Lemma 5.10, we get that for an arbitrary $A\in\bigvee_{i=0}^{n-1}f^{-i}(\mathcal{W}_{m})$ where $\mathcal{W}_{m}\coloneqq\{U\cap X:U\in\mathcal{U}_{m}\}$, the number of sets in the open cover $\bigvee_{j=0}^{n-1}f^{-j}(\mathcal{W}_{l})$ needed to cover $A$ can be bounded from above, and such an upper bound leads to the conclusion that $h^{*}(f)=0$. To prove Theorem B, given a sufficiently large $m\in\mathbb{N}$ we construct recursively a sequence $\{A_{n}\}_{n\in\mathbb{N}}$ of subsets of $X$ such that each $A_{n}$ is contained in an element of $\bigcap_{i=0}^{n}f^{-i}(U_{i}^{m}\cap X)$. We prove by induction that for each sufficiently large $l\in\mathbb{N}$, the number of elements in $\bigvee_{j=0}^{n-1}f^{-j}(\mathcal{W}_{l})$ needed to cover $A_{n}$ is bounded from below, and such lower bound leads to the conclusion that $h(f|\mathcal{W}_{m})>0$.
In Section 6, we show the existence of equilibrium states for a metric CXC system and a continuous potential, which is a consequence of the asymptotic $h$-expansiveness.
In Section 7, with some additional assumptions, we study a level-$2$ large deviation principle and an equidistribution result for iterated preimages.
In Subsection 7.1, we briefly review a level-$2$ large deviation principle in our context, and record in Theorem 7.1 a theorem of Y. Kifer [Ki90], reformulated by H. Comman and J. Rivera-Letelier [CRL11], on level-$2$ large deviation principles. This result will be applied later to our context.
In Subsection 7.2, we state the additional assumptions, which will be repeatedly referred to as the Additional Assumptions in Subsection 7.2.
Subsection 7.3 consists of the proof of Theorem D. We first recall the geometric coding tree and the semiconjugacy from a shift map to a CXC system. Then, with this idea, we generalize some characterization of topological pressure in Lemma 7.18. This characterization of topological pressure allows us to verify the conditions in Theorem 7.1, and thus apply Theorem 7.1 to show a level-$2$ large deviation principle for iterated preimages.
Subsection 7.4 includes a proof of Theorem E. In Corollary 7.19, we obtain a characterization of the measure-theoretic pressure in terms of the infimum of certain limits involving iterated preimages. Such a characterization is then used in the proof of the equidistribution results.
@@ 2. Notation
In this section, we set up some notations for this paper. Let $\mathbb{Z}$ be the set of integers, {1, 2, $\mathbb{N}\coloneqq\{1,\,2,\,3,\,\dots\}$ be the set of positive integers, and $\mathbb{N}_{0}\coloneqq\{0\}\cup\mathbb{N}$. The symbols $\log$ and $\exp$ denote the logarithm and exponential function to the base $e$, respectively. The symbol $\operatorname{card}(A)$ denotes the cardinality of a set $A$. For $x\in\mathbb{R}$, define $\lceil x\rceil\coloneqq\min\{n\in\mathbb{Z}:n\geq x\}$ and $\lfloor x\rfloor\coloneqq\max\{n\in\mathbb{Z}:n\leq x\}$.
@@
Let $C(X)$ be the space of continuous functions from $X$ to $\mathbb{R}$. For $\alpha\in(0,1]$, $C^{0,\alpha}(X)$ is the space of $\alpha$-Hölder continuous function from $(X,d)$ to $\mathbb{R}$. Let $\mathcal{P}(X)$ be the set of Borel probability measures on $X$. For $g:X\rightarrow X$, let $\mathcal{M}(g,X)$ be the set of $g$-invariant Borel probability measures on $X$. For a point $x\in X$, we denote the Dirac measure supported on $\{x\}$ by $\delta_{x}$.
@@ 3.1. Finite branched covering maps
We define here the notion of finite branched covering maps. FBC maps generalize some rational maps on the Riemann sphere, capturing their essential topological properties. Here we refer to [HP09], where the reader can find a more detailed introduction.
@@
Suppose that $X$ and $Y$ are locally compact Hausdorff spaces, and let $f\colon X\rightarrow Y$ be a finite-to-one continuous map. The map $f$ is a finite branched covering (abbr: FBC) map if $\deg(f)<+\infty$ and the following conditions are satisfied.
@@
We give some basic properties of FBC maps below. Further details can be found in Section 2.1 of [HP09].
The composition of two FBC maps is a FBC map. The degrees and local degrees multiply under composition. Condition (ii) in Definition 3.1 implies that if a sequence of points $\{x_{n}\}$ converges to $x_{0}$, then $\deg(f;x_{n})\leq\deg(f;x_{0})$ for all sufficiently large $n\in\mathbb{N}$. It follows that the branch set $B_{f}$ is closed. The set of principal values is $Y\setminus V_{f}$.
@@
Here we prove two technical properties of finite-to-one, open, closed, and surjective continuous maps.
@@
Proof.
Since $V$ is open in $Y$, $U$ is open in $X$. Then, by the assumption on $f$, it can be checked that $f|_{U}$ is finite-to-one, open, and surjective.
Next, we show that $f|_{U}$ is closed. Let $A\subseteq U$ be closed in $U$, then
$A=\overline{A}\cap U=\overline{A}\cap f^{-1}(V),$
where $\overline{A}$ is the closure of $A$ in $X$. It follows that $f(A)=f(\overline{A})\cap V$. Since $f$ is closed, $f(\overline{A})$ is closed in $Y$, then $f(A)$ is closed in $V$. Thus $f|_{U}$ is closed. ∎
@@
Proof.
Since $f$ is continuous, $f^{-1}(V)$ open in $X$. Then since $X$ is locally connected, for each $x\in U$, there exists a connected neighborhood $U_{x}$ of $x$ such that $U_{x}\subseteq f^{-1}(V)$, hence $U_{x}\subseteq U$. So $U$ is open in $X$. Then it can be checked that $f|_{U}$ is finite-to-one and open. To show that $f|_{U}$ is closed, let $A\subseteq U$ be closed in $U$, then
$A=\overline{A}\cap U=\overline{A}\cap f^{-1}(V),$
where $\overline{A}$ is the closure of $A$ in $X$. It follows that $f(A)=f(\overline{A})\cap V$. Since $f$ is closed, $f(A)$ is closed in $V$. Since $f(U)$ is open and closed in $V$, $f(U)=V$. Thus $f|_{U}$ is surjective. ∎

### v3

@@ Abstract.
Coarse expanding conformal systems were introduced by P. Haïssinsky and K. M. Pilgrim to study the essential dynamical properties of certain rational maps on the Riemann sphere in complex dynamics from the point of view of Sullivan’s dictionary.
In this paper, we prove that for a metric coarse expanding conformal system $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ with repellor $X$, the map $f|_{X}\colon X\rightarrow X$ is asymptotically $h$-expansive. Moreover, we show that $f|_{X}$ is not $h$-expansive if there exists at least one branch point in the repellor. We also prove that $f|_{X}$ is forward expansive when there is no branch point in the repellor.
Our study of expansion properties does not rely on symbolic codings or Markov partitions, and exploits directly the geometric data.
As a consequence of asymptotic $h$-expansiveness, for $f|_{X}$ and each real-valued continuous potential on $X$, there exists at least one equilibrium state. For such maps, if some additional assumptions are satisfied, we can furthermore establish a level-$2$ large deviation principle for iterated preimages, followed by an equidistribution result.
@@ 1.1. Background and motivation
In complex dynamics, the first-named author studied $h$-expansiveness and asymptotic $h$-expansiveness for expanding Thurston maps [Li15]. In this paper, we consider a more general context in complex dynamics. The dynamical systems that we study in this paper are called coarse expanding conformal (CXC) systems, introduced by P. Haïssinsky and K. M. Pilgrim [HP09]. CXC systems are a class of topological systems defined in an axiomatic way, generalizing a wide range of maps, including some non-uniformly expanding rational maps on the Riemann sphere $\widehat{\mathbb{C}}$ and expanding Thurston maps without periodic critical points studied by M. Bonk and D. Meyer [BM10, BM17] dating back to W. P. Thurston’s topological characterization of postcritially finite rational maps. Other examples of CXC systems include
@@
    Lattès type uniformly quasiregular maps on Riemannian manifolds;
@@
    smooth expanding partial self-covering maps of Riemannian manifolds. More precisely, let $\mathfrak{X}_{0}$ be a connected complete Riemannian manifold and $\mathfrak{X}_{1}\subseteq\mathfrak{X}_{0}$ an open submanifold with finitely many connected components which is compactly contained in $\mathfrak{X}_{0}$, then an expanding $C^{1}$ covering map $f\colon\mathfrak{X}_{0}\rightarrow\mathfrak{X}_{1}$ is CXC.
We refer the reader to [HP09, Sections 2.3, 4.2, and 4.4], and [CJY94, CT11, HP10, JZ09] for more information regarding examples listed above.
@@
In this paper we establish the asymptotic $h$-expansiveness of each metric CXC system $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ with repellor $X$, more precisely, of the map $f|_{X}$. We also study expansiveness and $h$-expansiveness of metric CXC systems. The motivation to investigate asymptotic $h$-expansiveness comes from the theory of thermodynamic formalism. Investigation on concepts such as topological entropy, measure-theoretic entropy, topological pressure, and measure-theoretic pressure is a main theme in thermodynamic formalism. The maximizing measures of measure-theoretic entropy and measure-theoretic pressure are known as measures of maximal entropy and equilibrium states, respectively. The existence, uniqueness, and various other properties of equilibrium states have been studied in many contexts (see for example, [Bo75, Ru89, Pr90, KH95, MauU03, BS03, Ol03, Yu03, PU10, MayU10]).
M. Misiurewicz showed that asymptotic $h$-expansiveness of a map $f$ guarantees that the measure-theoretic entropy $\mu\mapsto h_{\mu}(f)$ is upper semi-continuous [Mi76] with respect to the weak topology. It implies that for each real-valued continuous function $\phi$, the measure-theoretic pressure $\mu\mapsto P_{\mu}(f,\phi)$ is upper semi-continuous, which guarantees the existence of equilibrium states.
@@ 1.2. Main results
Then, we study forward expansiveness and $h$-expansiveness of metric CXC systems.
@@
Let $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ be a metric CXC system. Then, the following statements are equivalent:
  • (i)
    There is no branch point in $X$.
  • (ii)
    $f|_{X}$ is forward expansive.
  • (iii)
    $f|_{X}$ is $h$-expansive.
@@
Let $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ be a metric CXC system with repellor $X$. Then, for each real-valued continuous function $\phi\colon X\rightarrow\mathbb{R}$, there exists at least one equilibrium state for the map $f|_{X}$ and potential $\phi$. In particular, there exists at least one measure of maximal entropy for $f|_{X}$.
Applying some additional assumptions (see Subsection 7.2), we furthermore establish a level-$2$ large deviation principle (see Subsection 7.1 for the definition) for metric CXC systems. We refer the reader to Section 2 for the notations in the statement of the following theorem.
@@
satisfies a level-$2$ large deviation principle with rate function $I^{\phi}$, and it converges in the weak topology to the Dirac measure supported on the unique equilibrium state for $f$ and potential $\phi$. Furthermore, for each convex open subset $\mathfrak{G}$ of $\mathcal{P}(X)$ containing some invariant measure, we have
@@
Under proper assumptions, there exists a unique equilibrium state for the map $f|_{X}$ and each Hölder continuous potential due to [DPTUZ21, Theorem 1.1]. A consequence of the level-$2$ large deviation principle is the equidistribution of iterated preimages with respect to the equilibrium state (see Corollary 7.19).
Remark.
As shown in the recent work [LS24] of the first-named author and Xianghui Shi, for expanding Thurston maps, the absence of periodic critical points may not be essential for level-2 large deviation principles, even though the proof in the general case is much more involved relying on the construction of so-called subsystems. Subsystems rely on Markov partitions for expanding Thurston maps. It is therefore also interesting to investigate level-2 large deviation principles as in Theorem D without assuming the [Deg] axiom on $f$ and without using Markov partitions.
Our results in this paper partially extend results in [HP09, BM10, BM17, Li15, DPTUZ21].
Note that each expanding Thurston map $f\colon S^{2}\rightarrow S^{2}$ without periodic critical points is a metric CXC system (if we equip the sphere with a visual metric) with all branch points in $X$. So Theorems A and B strengthen Theorem 1.1 in [Li15] in the case where $f$ has no periodic critical points. In contrast, if an expanding Thurston map has a periodic critical point, then it is not a metric CXC system (or even a topological CXC system, see Subsection 3.2), and is known to be not asymptotically $h$-expansive (see [Li15, Theorem 1.1]). In the special case where the potential vanishes everywhere, the existence and uniqueness of the measure of maximal entropy and the corresponding equidistribution results were obtained by M. Bonk and D. Meyer [BM10, BM17] for expanding Thurston maps, and by P. Haïssinsky and K. M. Pilgrim [HP09] for CXC systems.
One major difference between a Thurston map and a CXC system is that the former is defined on a topological $2$-sphere, while the latter considers more general topological spaces, and that the former requires the set of forward iterated images of critical points to be finite, while the latter does not require such a strong assumption. This so-called postcritical finiteness assumption is crucial in the work of M. Bonk and D. Meyer [BM10, BM17] for the construction of certain forward invariant Jordan curves that induce Markov partitions for (sufficiently high iterations of) an expanding Thurston map (see also the work of J. W. Cannon, W. J. Floyd, and W. R. Parry [CFP07]). The proof of [Li15, Theorem 1.1] depends heavily on the combinatorial structures induced by the forward invariant Jordan curves and the corresponding Markov partitions, which are not available in the general setting of CXC systems. Our strategy in this paper builds upon the construction of bouquets (see Definition 5.8) from the more flexible good open covers of P. Haïssinsky and K. M. Pilgrim [HP09].
The main technical device $E_{m}(U_{0},\,\dots,\,U_{n-1};U_{n})$ (see (5.11)) is now defined using good open sets $U_{0},\,\dots,\,U_{n}\in\mathcal{U}_{m}$, which differs from similar constructions in [Li15] relying on Markov partitions. The lack of Markov partitions results in significant technical difficulties. The combinatorial data of expanding Thurston maps plays an important role in [Li15]. In [Li15], the combinatorial structure of so-called flowers, the bi-Lipschitz property with respect to visual metrics of an expanding Thurston map, and the finiteness of the postcritical set, etc., are widely used. On the other hand, in the current paper, we exploit directly the geometric data of CXC systems, for example, the roundness distortion properties and local comparability of the good open sets as well as the so-called bouquets.
We believe that our geometric approach, which does not rely on symbolic codings or Markov partitions, can be applied in other situations and inspire further research.
1.3. Structure of the paper
We will now give a brief description of the structure of this paper.
@@ 1.3. Structure of the paper
In Section 5, we investigate the weak expansion properties of metric CXC systems. We first introduce some basic concepts in Subsection 5.1. We review the notion of topological conditional entropy $h(g|\eta)$ of a continuous map $g\colon Z\rightarrow Z$ given an open cover $\eta$ of $Z$, and the notion of topological tail entropy $h^{*}(g)$. We then define $h$-expansiveness and asymptotic $h$-expansiveness using these notions. In Subsection 5.2, we prove several lemmas that will be used in the proof of the asymptotic $h$-expansiveness. Subsection 5.3 consists of the proofs of Theorems A, Theorem B.
In Section 6, we show the existence of equilibrium states for a metric CXC system and a continuous potential, which is a consequence of the asymptotic $h$-expansiveness. This section contains the proof of Theorem C.
In Section 7, with some additional assumptions, we study a level-$2$ large deviation principle and an equidistribution result for iterated preimages. In Subsection 7.1, we briefly review a level-$2$ large deviation principle in our context, and record in Theorem 7.1 a theorem of Y. Kifer [Ki90], reformulated by H. Comman and J. Rivera-Letelier [CRL11], on level-$2$ large deviation principles. This result will be applied later to our context. In Subsection 7.2, we state the additional assumptions, which will be repeatedly referred to as the Additional Assumptions in Subsection 7.2. Subsection 7.3 consists of the proof of Theorem D.
Acknowledgments
The first-named author would like to thank Peter Haïssinsky and Juan Rivera-Letelier for insightful discussions. The authors would like to thank the anonymous referees for their valuable suggestions. The authors were partially supported by NSFC Nos. 12471083, 12101017, 12090010, 12090015, BJNSF No. 1214021, and Peking University Funding No. 7101303303, .
@@ 2. Notation
In this section, we set up some notations for this paper. Let $\mathbb{Z}$ be the set of integers, $\mathbb{N}\coloneqq\{1,\,2,\,3,\,\dots\}$ be the set of positive integers, and $\mathbb{N}_{0}\coloneqq\{0\}\cup\mathbb{N}$. The symbols $\log$ and $\exp$ denote the logarithm and exponential function to the base $e$, respectively. The symbol $\operatorname{card}(A)$ denotes the cardinality of a set $A$. For $x\in\mathbb{R}$, define $\lceil x\rceil\coloneqq\min\{n\in\mathbb{Z}:n\geqslant x\}$ and $\lfloor x\rfloor\coloneqq\max\{n\in\mathbb{Z}:n\leqslant x\}$.
@@
Let $C(X)$ be the space of continuous functions from $X$ to $\mathbb{R}$. For $\alpha\in(0,1]$, $C^{0,\alpha}(X)$ is the space of $\alpha$-Hölder continuous function from $(X,d)$ to $\mathbb{R}$. Let $\mathcal{P}(X)$ be the set of Borel probability measures on $X$. For $g:X\rightarrow X$, let $\mathcal{M}(X,g)$ be the set of $g$-invariant Borel probability measures on $X$. For a point $x\in X$, we denote the Dirac measure supported on $\{x\}$ by $\delta_{x}$.
@@ 3.1. Finite branched covering maps
We define here the notion of finite branched covering maps. FBC maps generalize some rational maps on the Riemann sphere, capturing their essential topological properties. Here we refer to [HP09, Section 2.1], where the reader can find a more detailed introduction.
@@
Suppose that $X$ and $Y$ are locally compact Hausdorff spaces, and let $f\colon X\rightarrow Y$ be a finite-to-one continuous map. The map $f$ is a finite branched covering (abbr: FBC) map if $\deg(f)<+\infty$ and the following conditions are satisfied:
@@
We give some basic properties of FBC maps below. Further details can be found in [HP09, Section 2.1].
The composition of two FBC maps is a FBC map. The degrees and local degrees multiply under composition. Condition (ii) in Definition 3.1 implies that if a sequence of points $\{x_{n}\}$ converges to $x_{0}$, then $\deg(f;x_{n})\leqslant\deg(f;x_{0})$ for all sufficiently large $n\in\mathbb{N}$. It follows that the branch set $B_{f}$ is closed. The set of principal values is $Y\setminus V_{f}$.
@@
Here we give two technical properties of finite-to-one, open, closed, and surjective continuous maps. We skip the proofs since they are standard.